PLEASE COMPLETE ALL PARTS AND SHOW WORK

A local college reported that the average SAT score of its students is 1100 with a standard deviation of 45. Dr. Smith drew a random sample of 25 students from this college and found that her sample mean was 1082.

What is the z statistic for this sample? Show your work.

Is Dr. Smith’s sample representative of the students at this college? Explain your reasoning.
Answer #1

Z stat = ( xbar - mu )/ [ sigma ÷ sqrt ( n ) ]

Z stat = ( 1082 - 1100 ) / [ 45 ÷ sqrt( 25 ) ] = -2

Ho : mu = 1100

Ha : mu not equal 1100

Z test = - 2 < 1.96 = level of significance

Ho is not rejected

Yes it is representative of the students at this college